236.7 miles in 4.5 days
how many days for 1 mile?
331.5 pages in 8.5 weeks
how many weeks for 1 page?

Answer by checkley75(3666)

1/236.7=x/4.5 cross multiply
236.7x=4.5
x=4.5/236.7
x=.019 days for 1 mile.
-------------------------------------------
1/331.5=x/8.5 cross multiply
331.5x=8.5
x=8.5/331.5
x=.02564 weeks for 1 page